# Dual Conditional Diffusion Models for Sequential Recommendation

### v1

@@ 5. Experiment
$\bullet$ RQ1: How does the performance of DCRec  compare with other baselines across the datasets in different experiment settings?
$\bullet$ RQ2: What is the impact of different components (e.g., reconstruction loss, ranking loss, regularization loss, and DCDT  modules) within the DCRec  on overall performance?
$\bullet$ RQ3: How does DCRec  perform in terms of inference efficiency?
@@ 5.1. Experimental Settings
$\bullet$ Datasets. We evaluate our proposed DCRec  on three publicly available datasets: Amazon 5-core Beauty, Amazon 5-core Toys, and Yelp, following the experimental setup of previous works (Zhou et al., 2020a; Chen et al., 2022a; Li et al., 2024b). Detailed descriptions of these datasets can be found in Appendix B due to space constraints.
$\bullet$ Baselines. To evaluate the performance of our model, we select various baselines covering three types of mainstream methods in SR tasks, e.g., representative SR methods, contrastive learning SR methods, and diffusion-based SR methods. The detailed information is listed in the Appendix C.
$\bullet$ Implementation Details and Parameter Settings. To ensure fair comparisons, all the baselines and ours are implemented based on the popular recommendation framework RecBole (Zhao et al., 2021) under the same settings and evaluated consistently. The detailed information is listed in the Appendix F.
$\bullet$ Evaluation Metrics. We use two commonly used metrics including Normalized Discounted Cumulative Gain @K (NDCG@K), Hit Ratio @K (HR@K) to evaluate the performance of all compared methods. For all the baseline models, we generate the ranking list of items for each testing interaction. Both of them are applied with $K$ chosen from {5, 10}, evaluating the ranking results using all items in the dataset as candidates. Meanwhile, we follow the setting in (Wang et al., 2019b), using the paired t-test with p¡0.05 for the significance test.
5.2. Overall Performance (RQ1)
@@ 5.3. In-depth Analysis (RQ2)
The results compare four configurations: single attention (Single-attn) with and without Conditional Layer Normalization (CondLN) and dual attention (Dual-attn) with and without CondLN. Single-attn uses only a self-attention block where the noised sequence, diffusion step, and history embeddings are directly combined. In contrast, Dual-attn integrates both self-attention and cross-attention layers to capture interactions between the noised sequence and user history. CondLN conditions the LN process on the noised sequence embedding $S_{t}$, diffusion time embedding $t$ adapting to personalized user behaviors. The results show that both CondLN and dual attention effectively boost performance with highest HR@5 performance when combining them. Meanwhile, the absence of any of these modules results in a decline in performance.
The effectiveness of DCRec. To verify the rationality of incorporating both implicit conditional and explicit conditional effectiveness in DCRec, we choose implicit conditional components only, explicit conditional components only (broadcasting the target item to the same length as historical interactions), and the original dual conditional diffusion transformer. As shown in Table 2, DCRec  consistently outperforms both ICDM and ECDM, demonstrating that jointly modeling implicit and explicit historical conditions is more effective than relying on either one individually.
@@
For a more nuanced understanding, we select certain sensitive hyper-parameter $\lambda$, and explore three different $\lambda$ settings as shown in Figure 6.
Figure 6. Figure (a) depicts three types of $\lambda$ and Figure (b) reports the corresponding HR@5 results on Amazon Beauty and Toys datasets. The $\lambda_{smooth}$ reports superior performance.
The loss balance factor $\lambda(\cdot)$ influences DCRec  performance by mediating the focus between recommendation and reconstruction tasks. While an increase in $\lambda$ initially bolsters performance by prioritizing embedding denoising, too high a value risks neglecting the core recommendation task, undermining overall performance. $\lambda_{fix}$ denotes experimental constant, $\lambda_{decline}$ denotes handcrafted function, $\lambda_{smooth}$ denotes training epoch related function, respectively. We can find that $\lambda_{smooth}$ can achieve the best performance. Due to the page limit, the details of $\lambda$ and other hyper-parameters please refer to Appendix E.
5.4. Model Efficiency Analysis (RQ3)
@@
In this section, we investigate the impact of inference step $T$ to answer RQ3. We vary diffusion steps $T$ from 0 to $T$ in both Beauty and Toys datasets during inference and show the results in Figure 7. Moreover, we conducted a time comparison with DiffuRec (Li et al., 2024b), the most representative method in the baseline. To the fair comparison, we set 30 diffusion steps for training our DCRec  aligning to the DiffuRec as shown in Figure 8.
Firstly, We can find the best inference steps are vary in different datasets. This implies that the optimal diffusion step for each dataset is related to their item count. Then we observe that our DCDT ($f_{\theta}$) is optimized powerful enough to fulfill Equation 36 at an early $\hat{t}$-th stage, which significantly reduces time overhead. Compared to DiffuRec, our DCRec’s performance is more stable under different inference steps $T$ settings. In conclusion, our DCRec  can reach superior model performance than DiffuRec while maintaining lower inference time cost under different inference step settings. This is because our improvement enables the dual conditional mechanism, resulting in both accurate and efficient inference.
@@ 6. Conclusion
In this paper, we proposed the DCRec  to address key limitations in existing diffusion-based sequential recommender systems. By introducing a Markov chain to model item transitions and a dual conditional module to integrate contextual information, our approach improves the consistency of item selection and enhances the generation of target item embedding. Extensive experiments on public datasets validate the effectiveness of DCRec, demonstrating its superiority over state-of-the-art methods in both recommendation accuracy and computational efficiency.
Appendix A Algorithm of DCRec  Training and Inference
Algorithm 1 DCRec  Training
2:repeat
3:     Sample ${H_{0}}$ from $\{H_{0}\}$ and corresponding $z$ from $\mathbb{Z}$
4:     $S_{0}$leftarrow Concatenate $H_{0}$ and $z$
5:     Sample $t\sim\mathcal{U}(1,T)$
6:     $S_{t}$leftarrow Adding noise to $S_{0}$ according to Equation 23
7:     Obtain predicted $\hat{S}_{0}\leftarrow f_{\theta}(S_{t},t,H_{0})$
8:     Extract predicted $\hat{e}_{0}$ from $\hat{S}_{0}$
9:     Obtain predicted $\hat{z}\leftarrow p_{\phi}(\hat{z}|\hat{e}_{0})$
10:     $\mathcal{L}_{T}$leftarrow $\|\hat{S}_{0}\|^{2}$
11:     $\mathcal{L}_{t}$leftarrow $\|\hat{e}_{0}-e_{0}\|^{2}$
12:     $\mathcal{L}_{z}$leftarrow $D_{\text{KL}}(q_{\phi}(z)||p_{\phi}(\hat{z}))$
14:until converged
15:optimized $\theta$, $\phi$.
Algorithm 2 DCRec  Inference
2:Uniformly sample $T^{\prime}$ steps from $1$ to $T$
3:Sample $e_{T}~{}\sim\mathcal{N}(\bm{0},\bm{I})$
4:Sample $H_{T}~{}\sim\mathcal{N}(\bm{0},\bm{I})$
5:Implicit conditioning $H_{T}\leftarrow H_{T}+\delta H_{0}$
6:$S_{T}\leftarrow$ Concatenate $H_{T}$ and $e_{T}$
7:for $t\leftarrow T^{\prime}\dots 1$  do
8:     Obtain predicted ${S}_{0}\leftarrow f_{\theta}(S_{t},t,H_{0})$
9:     Transit to $S_{t-1}\leftarrow q(S_{t-1}|{S}_{0})$
10:Extract $e_{0}$ from $S_{0}$
11:obtain $z\leftarrow p_{\phi}(z|e_{0})$
12:$z$
Appendix B Dataset Preparation
We evaluate our proposed DCRec  on three publicly accessible datasets in different experiment settings. Table 3 provides a summary of the statistics for each dataset.
  • Amazon Beauty & Amazon Toys: We choose three representative sub-datasets from Amazon datasets: Beauty and Toys and keep the ‘5-core’ datasets (Zhou et al., 2020a) (Rendle et al., 2010)(Chen et al., 2022a), which filter out user-item interaction sequences with length less than 5.
For reproducibility, we follow the commonly used benchmark setting of RecBole (Zhao et al., 2021) to set up our experiments. Specifically, for each user, we first discard duplicated interaction sequences and sort the items in each user’s sequence chronologically by their timestamp. Furthermore, the maximum length of interaction sequences is set to 50. If there are more than 50 interactions in a sequence, we adopt the latest 50 interactions. If the number of interactions in a sequence is less than 50, we make it up to 50 by padding virtual items with the ID of 0. Following the common practice in sequential recommendations, we leave the interactions happening at the latest time as the test dataset and the interactions at the second latest time as the validation dataset (Chen et al., 2022a; Zhou et al., 2020b).
Table 3. Statistics of the datasets after preprocessing.
Dataset Beauty Toys Yelp
# Users 22,363 19,412 30,499
# Items 12,101 11,924 20,068
# Avg. Actions / User 8.9 8.6 10.4
# Avg. Actions / Item 16.4 14.1 15.8
# Actions 198,502 167,597 317,182
Sparsity 99.93% 99.93% 99.95%
Appendix C Baselines for Comparison
To evaluate the performance of our model DCRec, we select various representative and/or state-of-the-art recommendation models as baselines.
  • Caser (Yuan et al., 2019): CNN-based approach for SR.
  • $S^{3}$ec (Zhou et al., 2020c): uses SSL to capture correlation-ship among items, sub-sequences, and associated attributes from the given user behavior sequence. Its modules for mining on attributes are removed because we don’t have attributes for items.
  • SdifRec (Xie et al., 2024): introduce the Schrödinger Bridge into diffusion-based SR, by replacing the Gaussian prior with the user’s current state.
  • DiffuRec (Li et al., 2024b): incorporates a diffusion model into the field of SR, leveraging a Transformer backbone to reconstruct target item representations based on the user’s historical interaction behaviors.
Appendix D Ablation Study for Partially Noised Input Sequence
In this paper, we concatenate the history embedding and target item embedding as a sequential input for our denoising model DCDT. As shown in Figure 2, we add noise to the whole input sequence. In this section, motivated by DiffuSeq (Gong et al., 2023) which is another diffusion model for Q&A task in NLP, we analyze the effectiveness of only adding noise to the target item. The results are shown in Figure 9. It seems that partially adding noise to the target item is worse than adding noise to the whole sequence. The possible reason is the unbalanced noised sequence length. DiffuSeq concatenates the question and answer as an input and only adds noise to the answer sequence. In that case, the question sequence and answer sequence have similar sequence lengths. However, in our SR task, the history sequence is nearly 50×50$\times$50 × longer than the target. Therefore, the strategy of partially noised input is unfeasible for DCRec.
Figure 9. Performance comparison of history conditioning with noise and without noise.
Appendix E Details of Hyper-parameters Analysis
E.1. The Balance Factor $\lambda$
It’s worth to note that the optimization in Equations 31 and 24 is a complex and parameter-sensitive multi-objective task. In this section, we detailed analyze the impacts of three types of different $\lambda$ settings, including fix factor $\lambda_{\text{fix}}$, decline factor $\lambda_{\text{decline}}$ and smoothly decline factor $\lambda_{\text{smooth}}$. We formulate them as follows:
$\lambda_{\text{fix}}&=c\\ \lambda_{\text{decline}}&=\text{Max}-g_{d}(\text{ep})+\text{Min}\\ \lambda_{\text{smooth}}&=\text{Max}*g_{s}(\text{ep})+\text{Min}\\ \end{aligned}\right.$
where $c$ is a constant and ep is the current training epoch. Max and Min denote the upper bound and the lower bound of the decline factor $\lambda_{\text{decline}}$ and $\lambda_{\text{smooth}}$. $g_{d}(\text{ep})$ is a decline linear function associated with the current training epoch, while $g_{s}(\text{ep})$ is another decline exponential function also associated with the epoch. In Table 4 and Figure 10, we illustrate several example settings of $\lambda$ in our experiments and corresponding results are shown in Table 5. We note that those declining settings dramatically improve the model performance compared to the fixed ones. The $\lambda_{\text{smooth}}$ settings are feasible for Toys and Beauty datasets, while $\lambda_{\text{decline}}$ settings are more suitable for the Yelp dataset.
Table 4. Detailed settings of different $\bm{\lambda}$.
Tag Type Function
$\lambda_{\text{f1}}$ $\lambda_{\text{fix}}$ 0.2
$\lambda_{\text{f2}}$ $\lambda_{\text{fix}}$ 0.1
$\lambda_{\text{f3}}$ $\lambda_{\text{fix}}$ 0.003
$\lambda_{\text{d1}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/20+3\times 20^{-3}&0<\text{ep}<20\\ 3\times 10^{-3}&\text{ep}>=20\\ \end{cases}$
$\lambda_{\text{d2}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/30+3\times 10^{-3}&0<\text{ep}<30\\ 3\times 10^{-3}&\text{ep}>=30\\ \end{cases}$
$\lambda_{\text{d3}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/40+3\times 10^{-3}&0<\text{ep}<40\\ 3\times 10^{-3}&\text{ep}>=40\\ \end{cases}$
$\lambda_{\text{d4}}$ $\lambda_{\text{decline}}$$\begin{cases}0.2-0.2*\text{ep}/80+3\times 10^{-3}&0<\text{ep}<80\\ 3\times 10^{-3}&\text{ep}>=80\\ \end{cases}$
$\lambda_{\text{s1}}$ $\lambda_{\text{smooth}}$$\begin{cases}2*(1-\text{ep}/150)^{6}+3\times 10^{-3}&0<\text{ep}<150\\ 3\times 10^{-3}&\text{ep}>=150\\ \end{cases}$
$\lambda_{\text{s2}}$ $\lambda_{\text{smooth}}$$\begin{cases}2*(1-\text{ep}/150)^{9}+3\times 10^{-3}&0<\text{ep}<150\\ 3\times 10^{-3}&\text{ep}>=150\\ \end{cases}$
$\lambda_{\text{s3}}$ $\lambda_{\text{smooth}}$$\begin{cases}0.05*(1-\text{ep}/150)^{10}+3\times 10^{-3}&0<\text{ep}<150\\ 3\times 10^{-3}&\text{ep}>=150\\ \end{cases}$
Figure 10. Illustrations of different $\bm{\lambda}$.
Table 5. The comparison performance of different $\bm{\lambda}$ settings. Bold scores represent the highest results.
Dataset Metric $\lambda_{\text{fix}}$ $\lambda_{\text{decline}}$ $\lambda_{\text{smooth}}$
$\lambda_{f1}$ $\lambda_{f2}$ $\lambda_{f3}$ $\lambda_{d1}$ $\lambda_{d2}$ $\lambda_{d3}$ $\lambda_{d4}$ $\lambda_{s1}$ $\lambda_{s2}$ $\lambda_{s3}$
Beauty HR@5 0.0453 0.0507 0.0493 0.0320 0.0426 0.0479 0.0584 0.0617 0.0630 0.0564
HR@10 0.0643 0.0704 0.0669 0.0471 0.0636 0.0688 0.0844 0.0855 0.0859 0.0795
NDCG@5 0.0318 0.0366 0.0357 0.0224 0.0298 0.0340 0.0414 0.0435 0.0449 0.0399
NDCG@10 0.0379 0.0429 0.0413 0.0272 0.0365 0.4065 0.0498 0.0514 0.0522 0.0473
Toys HR@5 0.0470 0.0582 0.0458 0.0113 0.0284 0.0396 0.0676 0.0684 0.0690 0.0594
HR@10 0.0627 0.0787 0.0648 0.0195 0.0418 0.0552 0.0889 0.0901 0.0903 0.0804
NDCG@5 0.0352 0.0436 0.0329 0.0083 0.0203 0.0287 0.0503 0.0510 0.0518 0.0447
NDCG@10 0.0402 0.0502 0.0390 0.0109 0.0244 0.0337 0.0571 0.0579 0.0587 0.0514
Yelp HR@5 0.0241 0.0269 0.0247 0.0344 0.0405 0.0396 0.0251 0.0327 0.0289 0.0398
HR@10 0.0363 0.0411 0.0424 0.0547 0.0611 0.0605 0.0174 0.0514 0.0192 0.0591
NDCG@5 0.0169 0.0179 0.0163 0.0232 0.0272 0.0266 0.0388 0.0223 0.0443 0.0270
NDCG@10 0.0207 0.0225 0.0220 0.0297 0.0338 0.0333 0.0218 0.0282 0.0241 0.0333
E.2. Other Hyper-parameters Analysis
Figure 11. Hyperparameters analysis of our DCRec  on Beauty dataset.
We report the experimental results for key hyperparameters on the Beauty dataset, including the noise schedule $\beta$, number of diffusion steps $T$, and scaling factor $\delta$. In detail, $\beta$ is tuned within $\{0.01,0.02,0.04,0.06,0.08,0.1\}$, diffusion steps $T$ is tuned within $\{5,10,15,20,30,50,75,100,150,200\}$ and scaling factor $\delta$ is tuned within $\{0.1,0.2,0.5,0.75,1\}$. As shown in Figure 11, we can find for the noise schedule $\beta$, performance remains stable for values up to 0.04 but sharply declines at 0.06, indicating that excessive noise negatively impacts the model. The number of training diffusion steps $T$ improves performance up to 20 steps, after which it plateaus, suggesting that 20 steps are sufficient. The scaling factor $\delta$ performs best in the range of 0.2 to 0.5, with both smaller and larger values resulting in decreased performance.
Appendix F Implementation Details
We implement our model and baselines using Pytorch (Paszke et al., 2019) and conduct our experiments on an NVIDIA RTX 3090 GPU with 24G memory. For a fair comparison, we carefully tuned the model parameters of all the baselines on the validation set. Then we choose the model parameters that achieve the best performance on the validation set to compare. For our DCRec, based on the hyperparameter analysis, we set the batch size to 512, the item embedding dimension to 128, learning rate of $34\times 1e-$ 3, maximum sequence length to 50, $\beta$ to 0.02, the dropout rate to 0.1, the embedding dropout to 0.4, the hidden factor to 128, $\delta$ to 0.1, the number of transformer blocks to 4, and the number of diffusion steps $T$ to 50. For $\lambda$, we adopt $\lambda_{smooth}=\lambda_{s2}$ for the Beauty and Toys datasets, while for the Yelp dataset, we use $\lambda_{decline}=\lambda_{d2}$, as determined by the results in Table 5 result. We choose Adam (Kingma and Ba, 2015) as our optimization strategy and use early stopping with a patience of 5 to prevent overfitting.

### v2

@@ 4. Experiments
  • RQ1: How does the performance of DCRec  compare with other baselines across the datasets in different experiment settings?
  • RQ2: What is the impact of different components (e.g., concatenated input, reconstruction loss, ranking loss, regularization loss, and DCDT  modules) within the DCRec  on overall performance?
  • RQ3: What is the different impact of implicit, and explicit conditioning in DCRec?
  • RQ4: How does DCRec  perform in terms of inference efficiency?
  • RQ5: How does the noise concatenated input effect the diffusion model performance?
@@ 4.1. Experimental Settings
4.1.1. Datasets.
We evaluate our proposed DCRec  on three publicly available datasets: Amazon 5-core Beauty, Amazon 5-core Toys, and Yelp, following the experimental setup of previous works (Zhou et al., 2020b; Chen et al., 2022a; Li et al., 2024b). We evaluate our proposed DCRec  on three publicly accessible datasets in different experiment settings. Table 1 provides a summary of the statistics for each dataset.
  • Amazon Beauty & Amazon Toys: We choose three representative sub-datasets from Amazon datasets: Beauty and Toys and keep the ‘5-core’ datasets (Zhou et al., 2020b) (Rendle et al., 2010)(Chen et al., 2022a), which filter out user-item interaction sequences with length less than 5.
For reproducibility, we follow the commonly used benchmark setting of RecBole (Zhao et al., 2021) to set up our experiments. For each user, we first discard duplicated interaction sequences and sort the items in each user’s sequence chronologically by their timestamp. The maximum length of interaction sequences is set to 50. If there are more than 50 interactions in a sequence, we adopt the latest 50 interactions. If the number of interactions in a sequence is less than 50, we make it up to 50 by padding virtual items with the ID of 0. Following the common practice in SR, we leave the interactions happening at the latest time as the test dataset and the interactions at the second latest time as the validation dataset (Chen et al., 2022a; Zhou et al., 2020c).
Table 1. Statistics of the datasets after preprocessing.
Dataset Beauty Toys Yelp
# Users 22,363 19,412 30,499
# Items 12,101 11,924 20,068
# Avg. Actions / User 8.9 8.6 10.4
# Avg. Actions / Item 16.4 14.1 15.8
# Actions 198,502 167,597 317,182
Sparsity 99.93% 99.93% 99.95%
@@
4.1.2. Baselines.
To evaluate the performance of our model, we select various baselines covering three types of mainstream methods in SR tasks, e.g., representative SR methods, contrastive learning SR methods, and diffusion-based SR methods. To evaluate the performance of our model DCRec, we select various representative and/or state-of-the-art recommendation models as baselines.
  • Caser (Yuan et al., 2019): CNN-based approach for SR.
  • S3Rec (Zhou et al., 2020a): uses SSL to capture correlation-ship among items, sub-sequences, and associated attributes from the given user behavior sequence. Its modules for mining on attributes are removed because we don’t have attributes for items.
  • SdifRec (Xie et al., 2024): introduces the Schrödinger Bridge into diffusion-based SR, replacing the Gaussian prior with the user’s state.
  • DiffuRec (Li et al., 2024b): incorporates a diffusion model into the field of SR, using a Transformer backbone to reconstruct target item representations based on the user’s historical interaction behaviors.
4.1.3. Implementation Details and Parameter Settings.
To ensure fair comparisons, all the baselines and ours are implemented based on the popular recommendation framework RecBole (Zhao et al., 2021) under the same settings and evaluated consistently. We implement our model and baselines using Pytorch (Paszke et al., 2019) and conduct our experiments on an NVIDIA RTX 3090 GPU with 24G memory. For a fair comparison, we carefully tuned the model parameters of all the baselines on the validation set. Then we choose the model parameters that achieve the best performance on the validation set to compare. For our DCRec, based on the hyperparameter analysis, we set the batch size to 512, the item embedding dimension to 128, learning rate of $34\times 1e-$ 3, maximum sequence length to 50, $\beta$ to 0.02, the dropout rate to 0.1, the embedding dropout to 0.4, the hidden factor to 128, $\delta$ to 0.1, the number of transformer blocks to 4, and the number of diffusion steps $T$ to 50. For $\lambda$, we adopt $\lambda_{\text{smooth}}$ for Beauty, Toys and Yelp datasets. For detailed settings and further analysis of $\lambda$ , please refer to Section 4.3.2. Besides, we choose Adam (Kingma and Ba, 2015) as our optimization strategy and use early stopping with a patience of 5 to prevent overfitting.
4.1.4. Evaluation Metrics.
We use two commonly used metrics including Normalized Discounted Cumulative Gain @K (NDCG@K), Hit Ratio @K (HR@K) to evaluate the performance of all compared methods. For all the baseline models, we generate the ranking list of items for each testing interaction. Both of them are applied with $K$ chosen from {5, 10}, evaluating the ranking results using all items in the dataset as candidates. Meanwhile, we follow the setting in (Wang et al., 2019b), using the paired t-test with p¡0.05 for the significance test.
4.2. Overall Performance (RQ1)
@@ 4.3. In-depth Analysis (RQ2)
The results compare four configurations: single attention (Single-attn) with and without Conditional Layer Normalization (CondLN) and dual attention (Dual-attn) with and without CondLN. Single-attn uses only a self-attention block where the noisy sequence, diffusion step, and history embeddings are directly combined. In contrast, Dual-attn integrates both self-attention and cross-attention layers to capture interactions between the noisy sequence and user history. CondLN conditions the LN process on the noisy sequence embedding $S_{t}$, diffusion time embedding $t$ adapting to personalized user behaviors. The results show that both CondLN and dual attention effectively boost performance with highest HR@5 performance when combining them. Meanwhile, the absence of any of these modules results in a decline in performance.
The effectiveness of DCRec. To verify the rationality of incorporating both implicit conditional and explicit conditional effectiveness in DCRec, we choose implicit conditional components only, explicit conditional components only (broadcasting the target item to the same length as historical interactions), and the original dual conditional diffusion transformer. As shown in Table 3, DCRec  consistently outperforms both ICDM and ECDM, demonstrating that jointly modeling implicit and explicit historical conditions is more effective than relying on either one individually. Meanwhile, compared to implicit conditional diffusion baselines in Table 2, our ICDM has shown competitive performance on Beauty and Yelp, and better performance on Toys.
@@
The loss balance factor $\lambda(\cdot)$ influences DCRec  performance by mediating the focus between recommendation and reconstruction tasks. While an increase in $\lambda$ initially bolsters performance by prioritizing embedding denoising, too high a value risks neglecting the core recommendation task, undermining overall performance. For a more nuanced understanding, we explore three different $\lambda$ settings as shown in follows:
$\lambda_{\text{fix}}&=c,\\ \lambda_{\text{decline}}&=\text{Max}-g_{d}(\text{ep})+\text{Min},\\ \lambda_{\text{smooth}}&=\text{Max}*g_{s}(\text{ep})+\text{Min},\\ \end{aligned}\right.$
where, $\lambda_{fix}$ denotes an experimental constant $c$. $\lambda_{decline}$ and $\lambda_{smooth}$ denotes epoch-related linear function $g_{d}(\text{ep}$ and epoch-related power function $g_{s}(\text{ep})$, respectively, with upper bound Max and lower bound Min. We evaluate these three types of $\lambda$ with setting: $c=0.1$, $\{\text{Max},\text{Min}\}\!=\!\{0.2,0.003\}$, $g_{d}\!=\!(0.2-\text{ep}/400)$, and $g_{s}\!=\!(1-\text{ep}/150)^{9}$ as shown in Figure 6(a). The results in Figure 6(b) shows that $\lambda_{smooth}$ can achieve better performance.
Figure 6. (a) depicts three types of $\lambda$ and (b) reports the corresponding HR@5 results on Amazon Beauty and Toys datasets. The $\lambda_{smooth}$ reports superior performance.
4.4. Model Conditioning Analysis (RQ3)
We notice that the implicit components capture long-term behavioral patterns well, while the explicit components preserve temporal dynamics. To verify our claim, we provide a comparison experiment on the Amazon Beauty dataset. First, we divide the dataset into two subsets based on sequence length: one containing sequences with more than 10 interactions and the other with fewer than 10 interactions. Next, we train implicit-only, explicit-only, and DCRec models on the original dataset and evaluate their performance on the two subsets. The implicit-only and explicit-only models only using corresponding components in DCDT. The HR@5 results are shown in Table 4, illustrating that the implicit component captures long-term behavioral patterns well, while the explicit component preserves temporal dynamics.
Table 4. The effects of different conditioning approaches on different datasets with varying data length (HR@5).
Dataset Number of Sequence DCRec Implicit Explicit
Original 22,363 0.0630 0.0539 0.0585
Short Seq 17,240 0.0591 0.0488 0.0501
Long Seq 5,123 0.0639 0.0651 0.0635
4.5. Model Efficiency Analysis (RQ4)
@@
In this section, we analyze the network complexity of DCDT  and investigate the impact of inference step $T$ to answer RQ4. DCRec  has the same overall network complexity $\mathcal{O}(nd^{2}+dn^{2})$ as DreamRec and DiffuRec, where $n$ denotes the sequence length and $d$ is the embedding dimension. Besides, we vary diffusion steps from 0 to $T$ in both Beauty and Toys datasets during inference and show the results in Figure 7. We observe that the optimal inference steps vary across different datasets, suggesting that the ideal diffusion step for each dataset is influenced by its item count. Moreover, we conducted a time comparison with a strong baseline, DiffuRec (Li et al., 2024b). To the fair comparison, we set 30 diffusion steps for training our DCRec  aligning to the DiffuRec as shown in Figure 8. We observe that our DCDT ($f_{\theta}$) is optimized powerful enough to fulfill Equation 26 at an early $\bar{t}$-th stage, which significantly reduces time overhead. Compared to DiffuRec, our DCRec’s performance is more stable under different inference steps $T$ settings. In conclusion, our DCRec  shows superior model performance than DiffuRec while maintaining lower inference cost under different step settings. This is because our improvement enables the dual conditional mechanism, resulting in both accuracy and efficiency.
4.6. Ablation Study for Noise Adding (RQ5)
In this paper, we concatenate the history embedding and target item embedding as a sequential input for our denoising network DCDT. As shown in Figure 2, we add noise to the whole input sequence. In this section, motivated by DiffuSeq (Gong et al., 2023) which is another diffusion model for Q&A task in NLP, we analyze the effectiveness of only adding noise to the target item. The results are shown in Figure 9.
Partially adding noise to the target item proves less effective than adding noise to the whole sequence. The possible reason is the unbalanced noisy sequence length. DiffuSeq concatenates the question and answer as an input and only adds noise to the answer sequence. In that case, the question sequence and answer sequence have similar sequence lengths. However, in our SR task, the history sequence is nearly 50×50$\times$50 × longer than the target. Therefore, the strategy of partially noisy input is unfeasible for DCRec.
Figure 9. The impact of history conditioning with noise and without noise on Beauty and Toys datasets, reflecting by HR@5 and NDCG@5.
5. Related Work
5.1. Diffusion Models
Diffusion models, derived from non-equilibrium thermodynamics, have achieved notable success in areas like computer vision, sequence modeling, and audio processing (Yang et al., 2024b). Two primary paradigms exist: unconditional models and conditional models. Unconditional models, such as DDPM (Ho et al., 2020), aim to push the performance boundaries of generative models. Conditional models incorporate guidance from labels to control generated samples. SGGM (Yang et al., 2022) conditions on self-produced labels, while You et al. (You et al., 2023) demonstrate how few labels and pseudo-training enhance both large-scale diffusion models and semi-supervised learners. Dhariwal (Dhariwal and Nichol, 2021) uses classifier guidance to improve sample quality, and Ho and Salimans (Ho and Salimans, 2022b) combine conditional and unconditional scores to balance sample quality and diversity. Multi-modal guidance further enriches the diffusion process, as seen in DiffuSeq (Gong et al., 2023) for NLP tasks and SDEdit (Meng et al., 2022) for image translation, with latent diffusion models (LDM) (Rombach et al., 2022) offering flexible, unified semantics. unCLIP (Ramesh et al., 2022) and ConPreDiff (Yang et al., 2023a) integrate CLIP latents in text-to-image generation.
5.2. Diffusion for Sequential Recommendation
Diffusion models show promise in the sequential recommendation by using user behavior sequences as guidance in the denoising process (Xie et al., 2024; Yang et al., 2023b; Li et al., 2024b). DiffuRec (Li et al., 2024b) corrupts target item embeddings with Gaussian noise and uses a Transformer to iteratively reconstruct them. DreamRec (Yang et al., 2023b) generates the oracle next-item embeddings based on user preferences but struggles with scalability due to its lack of negative sampling. While promising to some extent, these implicit conditional methods may oversimplify complex user behaviors by compressing them into a single vector, potentially losing valuable sequential patterns and contextual information. Additionally, DCDR (Lin et al., 2024a) adopts a discrete diffusion framework for progressive re-ranking, generating a sequence rather than a specific target item.
@@ 6. Conclusion
In this paper, we proposed the DCRec  to address key limitations in existing diffusion-based sequential recommender systems. We first propose the Dual Conditional Diffusion Models for the sequential recommendation, introducing a novel diffusion framework for the SR. Based on this framework, we propose the DCDT  which can generate accurate target item embedding based on both implicit and explicit conditions. Besides, our framework simplifies the diffusion process by reducing some inference steps with high similarity. Extensive experiments on public datasets validate the effectiveness of DCRec, demonstrating its superiority over state-of-the-art methods in both recommendation accuracy and computational efficiency.